OpenStudy (anonymous):

if f(X)=x/3 and g(X)=1/x, find f(X)-g(X) a. x-1/3-x b. x^2-3/3x c. 1/3 d. 3-x^2/3x

OpenStudy (hoblos):

f(x)-g(x) = x/3 - 1/x = x²/3x - 3/3x = (x²-3)/3x

OpenStudy (anonymous):

true